# Searching for dark photons from dark-scalar decays at CEPC and FCC-ee

### v1

@@ I Introduction
One of the most profound mysteries in modern cosmology and particle physics is the nature of dark matter (DM). The existence of DM is strongly indicated by multiple astronomical observations, including gravitational lensing, the anisotropies of the cosmic microwave background, and the distribution of galaxies in large-scale surveys. Despite its overwhelming abundance, the fundamental nature of DM remains elusive. Various DM candidates have been proposed, ranging from weakly interacting massive particles and axion-like particles to dark gauge bosons, dark fermions, and other exotic particles predicted in extensions of the Standard Model (SM) Beacham et al. (2020); Agrawal et al. (2021); Antel et al. (2023).
In recent years, a class of models called “portal-physics” have been put under scrutiny by both theorists and experimentalists, considering both their simple constructions and strong physics motivations. They are supposed to be the portals that connect the visible sector (the SM spectrum) to a dark sector consisting of the DM, and often include dark photons Okun (1982); Galison and Manohar (1984); Holdom (1986); Boehm and Fayet (2004); Pospelov (2009); Curtin et al. (2015) or dark scalars O’Connell et al. (2007); Wells (2008); Bird et al. (2004); Pospelov et al. (2008); Krnjaic (2016); Boiarska et al. (2019). The dark photon $\gamma^{\prime}$ is a mediator of a hypothetical, broken dark $U(1)_{D}$ gauge symmetry that kinetically mixes with the SM photon, via the mixing term $\epsilon F_{\mu\nu}F^{\prime\mu\nu}$, where $F_{\mu\nu}$ and $F^{\prime}_{\mu\nu}$ are the field strength tensors of the SM photon and the dark photon, respectively, and $\epsilon$ is the kinetic mixing parameter. Through the kinetic mixing, the dark photon interacts with the SM particles with a coupling strength proportional to $\epsilon e$, where $e$ is the electromagnetic coupling constant.
@@
Furthermore, in some model constructions beyond the SM (BSM), such as the Hidden Abelian Higgs Model Wells (2008); Gopalakrishna et al. (2008), both the dark scalar and the dark photon are present where the dark photon acquires mass via the Higgs mechanism of the dark scalar. In this work, we focus on the low-energy phenomenology of this model and investigate, with the technique of Monte Carlo (MC) simulation, the search prospect of long-lived dark photons produced from dark-scalar decays at the next-generation electron-positron colliders, CEPC Dong et al. (2018); Abdallah et al. (2024); Adhya et al. (2025); Ai et al. (2025) and FCC-ee Abada et al. (2019). We restrict ourselves to the operation mode as the Higgs factories running at the center-of-mass (COM) energy $\sqrt{s}=240$ GeV, where the dark scalar can be singly produced via its mixing with $h$: $e^{+}e^{-}\to Z^{*}\to Z\,\phi$, and the dark photons are pair produced subsequently by the prompt decay of $\phi$: $\phi\to\gamma^{\prime}\gamma^{\prime}$. To our knowledge, searches for long-lived dark photons in this signal process have not been studied. For other phenomenological studies on theoretical scenarios involving both a dark Higgs boson and a dark photon, we refer to, e.g., Refs. Wells (2008); Ahlers et al. (2008); Gopalakrishna et al. (2008); Batell et al. (2009); Weihs and Zurita (2012); Davoudiasl et al. (2013); Curtin et al. (2014); Falkowski and Vega-Morales (2014); Curtin et al. (2015); Bakhet et al. (2015); Izaguirre and Stolarski (2018); Jodłowski et al. (2020); Araki et al. (2021); Foguel et al. (2022); Cheung et al. (2024); Araki et al. (2024).
@@
We confine ourselves to the mass ranges $m_{\phi}\in[25~\text{GeV},140~\text{GeV}]$ and $m_{\gamma^{\prime}}\in[2~\text{GeV},m_{\phi}/2]$. For these mass ranges, the current bounds on the mixing parameters $\epsilon$ and $\theta$ stem primarily from collider experiments. For the dark photon, the leading bounds on $\epsilon$ were obtained at the LHC collaborations CMS Sirunyan et al. (2020) and LHCb Aaij et al. (2020), as well as the $B$-factory experiment BaBar Lees et al. (2014), roughly all in the order of $10^{-3}$.
@@
For $m_{\phi}$ between 100 GeV and 1 TeV, constraints around $\theta\lesssim\mathcal{O}(10^{-1})$ were achieved by searches at CMS Khachatryan et al. (2015); Sirunyan et al. (2018a); Tumasyan et al. (2021) and ATLAS Aaboud et al. (2018); Aad et al. (2020, 2021, 2023). For $m_{\phi}$ between 10 GeV and 100 GeV, searches at CMS Sirunyan et al. (2018b), ATLAS Aad et al. (2022), LEP2 Barate et al. (2003), and L3 Acciarri et al. (1996) have placed upper bounds $\theta\lesssim\mathcal{O}(10^{-2})$. See Ref. Ferber et al. (2024) for a detailed summary of these existing bounds. In this work, we will simply assume $\theta=10^{-2}$ in numerical studies.
@@
For existing studies on the search prospect of dark photons and dark scalars at CEPC and FCC-ee, we refer to, for instance, Refs. He et al. (2018, 2017); Biswas et al. (2017); Park et al. (2023); Liu et al. (2017); İnan and Kisselev (2022); Jiang et al. (2020); Chang et al. (2018) and Refs. Liu et al. (2017); Chang et al. (2018); Cheung and Wang (2020); Alipour-Fard et al. (2019); Wang and Wang (2020); Ripellino et al. (2025); Bhattacherjee et al. (2025), respectively.
@@ II Theoretical model
We consider an extension of the SM that introduces an additional dark-sector gauge symmetry $U(1)_{D}$, under which all SM fields are uncharged. The new model includes both a dark photon $\gamma^{\prime}$ and a dark scalar $\phi$. After spontaneous symmetry breaking of the $U(1)_{D}$ via the vacuum expectation value (VEV) of the associated dark-scalar field, the dark photon acquires a mass $m_{\gamma^{\prime}}$ and an interaction between $\phi$ and two dark photons arises. The strength of this interaction scales with $m_{\gamma^{\prime}}$ and the $U(1)_{D}$ gauge coupling $g^{\prime}$. After the electroweak symmetry breaking, the dark photon will mix with the SM photon via a gauge kinetic-mixing term linking the SM hypercharge and the new $U(1)_{D}$ field strength tensors, $\epsilon F_{\mu\nu}F^{{}^{\prime}\mu\nu}$. In addition, the dark scalar $\phi$ mixes with the SM Higgs boson and thus couples to pairs of SM fermions, though the coupling strength is suppressed by the mixing parameter $\theta$. Thus, the low-energy phenomenological interaction Lagrangian employed in this study takes the form Araki et al. (2021):
@@
At the future Higgs factories, the SM Higgs bosons are dominantly produced via the Higgs-strahlung process, $e^{+}e^{-}\to Z\,h$. The dark scalar is produced in a very similar process, $e^{+}e^{-}\to Z\,\phi$, which is allowed as long as $\phi$ is light enough and it mixes with $h$. The cross section of the $e^{+}e^{-}\to Z\,\phi$ process is thus suppressed by $\theta^{2}$, and reads Djouadi (2008)
@@
Figure 1: The production cross section of $\phi$ via the process $e^{+}e^{-}\to Z\,\phi$ rescaled with $1/\theta^{2}$ as a function of the dark-scalar mass $m_{\phi}$, computed with MadGraph5 Alwall et al. (2014), at the COM energy $\sqrt{s}=240$ GeV.
In Fig. 1 we plot the scattering cross section $\sigma(e^{+}e^{-}\to Z\,\phi)$ rescaled with $1/\theta^{2}$ at the COM energy $\sqrt{s}=240$ GeV, as a function of $m_{\phi}$, computed with MadGraph5 3.4.2 Alwall et al. (2014). We find these results in agreement with Eq. (2).
As mentioned earlier, in this work, we focus on $m_{\phi}\sim 10\text{--}100$ GeV. With Eq. (1), the partial decay widths of the dark scalar into a pair of dark photons and a pair of lighter SM fermions can be calculated with Araki et al. (2021)
@@
The dark photon decays into a pair of SM charged fermions, with the partial decay widths Bauer et al. (2018); Batell et al. (2009); Araki et al. (2021); D’Onofrio et al. (2020); Fabbrichesi et al. (2020); Cheung et al. (2024)
@@ III The experiments and search strategies
Figure 5: A profile sketch of the inner tracker at CEPC and FCC-ee, extracted from Ref. Cheung and Wang (2020).
The main detectors of CEPC and FCC-ee share similar components and geometrical structures, while their exact dimensions differ to some extent. In particular, we will focus on the inner tracker (IT) component of the main detectors in our search strategies. We thus show a profile sketch of the IT at the CEPC or FCC-ee main detector in Fig. 5, reproduced from Ref. Cheung and Wang (2020). In Fig. 5, we denote an LLP $i$ with polar angle $\theta_{i}$ with a dashed line. $R_{I}$ ($R_{O}$) is the inner (outer) radius of the IT, and $L_{d}$ is the half length of the IT.
@@
Table 1: Parameters of the inner tracker at CEPC and FCC-ee, extracted from Ref. Dong et al. (2018) and Ref. Abada et al. (2019), respectively.
Further, we list in Table 1 the geometrical parameters of the inner trackers at CEPC Dong et al. (2018) and FCC-ee Abada et al. (2019). Here, $V$ denotes the volume.
@@ IV Simulation and calculation
We perform MC simulations with the tool Pythia8.3 Bierlich et al. (2022), scanning over $m_{\gamma^{\prime}}$ and $\tau_{\gamma^{\prime}}$ for representative values of $m_{\phi}$, in order to estimate the number of signal events $N_{S}$. Concretely, we select the following four values of $m_{\phi}$: 25, 50, 90, and 140 GeV, and for each value of $m_{\phi}$, we scan over $m_{\gamma^{\prime}}$ from 2 GeV to $m_{\phi}/2$. Specifically, at every scanned grid point of $(m_{\phi},m_{\gamma^{\prime}},\tau_{\gamma^{\prime}})$, we simulate ten thousand signal events. We then translate each combination of $(m_{\gamma^{\prime}},\tau_{\gamma^{\prime}})$ into $(m_{\gamma^{\prime}},\epsilon)$ with the help of Eq. (6) and the relation $\Gamma_{\text{total}}=\displaystyle\sum_{f}\Gamma_{\gamma^{\prime}\to f\bar{f}}=\hbar/\tau_{\gamma^{\prime}}$.
In Pythia8, we turn on the process HiggsSM:ffbar2HZ at a COM energy $\sqrt{s}=240$ GeV with back-to-back electron-positron collisions. By tuning the mass of the SM Higgs boson $h$ to match that of the dark scalar $\phi$, we mimic the kinematics of the signal process $e^{+}e^{-}\to Z\,\phi$. The dark scalar $\phi$ is then set to decay promptly and exclusively into a pair of long-lived dark photons. The kinematics of the dark photons in each signal event are then used as input for calculating the acceptance $\mathcal{A}_{i}$ of each search strategy; see Eq. (10).
@@
with $i=1,2$ for the two search strategies. $\sigma(e^{+}e^{-}\to Z\,\phi)$ is computed with MadGraph5; see Fig. 1 and the relevant discussion in Sec. II. Here, $\mathcal{L}_{\text{int.}}$ denotes the integrated luminosity, with $\mathcal{L}_{\text{int.}}=22$ ab-1 at CEPC Ai et al. (2025) and $\mathcal{L}_{\text{int.}}=5.6$ ab-1 at FCC-ee Abada et al. (2019).
For simplicity, we assume a detector efficiency of 100% in the calculation.
@@ V Sensitivity reach
Figure 6: Sensitivity results of the first (upper panel) and second (lower panel) strategies. Here, the COM energy is set at 240240 GeV and the scalar-Higgs mixing angle is fixed at $10^{-2}$. The gray-, green-, and pink-filled regions have been excluded by searches at LHCb Aaij et al. (2020), CMS Sirunyan et al. (2020), and BaBar Lees et al. (2014), respectively.
@@
The existing bounds from searches at LHCb Aaij et al. (2020), CMS Sirunyan et al. (2020), and BaBar Lees et al. (2014) are shown in the figures as color-filled areas.
@@
Finally, we comment that for $\theta$ values smaller than $10^{-2}$, worse sensitivity reach is expected, because fewer dark scalars and thus fewer dark photons would be produced. Quantitatively speaking, if $\theta$ is reduced by a factor of, say, 10, the total dark-photon production rates would decrease by 100. For the first search strategy, the lower sensitivity reach to $\epsilon$ would be weakened by a factor of 10 while that for the second search would be impaired by about a factor of $\sqrt{10}$.

### v2

@@ I Introduction
One of the most profound mysteries in modern cosmology and particle physics is the nature of dark matter (DM). The existence of DM is strongly indicated by multiple astronomical observations, including gravitational lensing, the anisotropies of the cosmic microwave background, and the distribution of galaxies in large-scale surveys. Despite its overwhelming abundance, the fundamental nature of DM remains elusive. Various DM candidates have been proposed, ranging from weakly interacting massive particles and axion-like particles to dark gauge bosons, dark fermions, and other exotic particles predicted in extensions of the Standard Model (SM) Beacham:2019nyx; Agrawal:2021dbo; Antel:2023hkf.
In recent years, a class of models called “portal-physics” have been put under scrutiny by both theorists and experimentalists, considering both their simple constructions and strong physics motivations. They are supposed to be the portals that connect the visible sector (the SM spectrum) to a dark sector consisting of the DM, and often include dark photons Okun:1982xi; Galison:1983pa; Holdom:1985ag; Boehm:2003hm; Pospelov:2008zw; Curtin:2014cca or dark scalars OConnell:2006rsp; Wells:2008xg; Bird:2004ts; Pospelov:2007mp; Krnjaic:2015mbs; Boiarska:2019jym. The dark photon $\gamma^{\prime}$ is a mediator of a hypothetical, broken dark $U(1)_{D}$ gauge symmetry that kinetically mixes with the SM photon, via the mixing term $\epsilon F_{\mu\nu}F^{\prime\mu\nu}$, where $F_{\mu\nu}$ and $F^{\prime}_{\mu\nu}$ are the field strength tensors of the SM photon and the dark photon, respectively, and $\epsilon$ is the kinetic mixing parameter. Through the kinetic mixing, the dark photon interacts with the SM particles with a coupling strength proportional to $\epsilon e$, where $e$ is the electromagnetic coupling constant.
@@
Furthermore, in some model constructions beyond the SM (BSM), such as the Hidden Abelian Higgs Model Wells:2008xg; Gopalakrishna:2008dv, both the dark scalar and the dark photon are present where the dark photon acquires mass via the Higgs mechanism of the dark scalar. In this work, we focus on the low-energy phenomenology of this model and investigate, with the technique of Monte Carlo (MC) simulation, the search prospect of long-lived dark photons produced from dark-scalar decays at the next-generation electron-positron colliders, CEPC CEPCStudyGroup:2018ghi; CEPCStudyGroup:2023quu; CEPCStudyGroup:2025kmw; Ai:2025cpj and FCC-ee FCC:2018evy. We restrict ourselves to the operation mode as the Higgs factories running at the center-of-mass (COM) energy $\sqrt{s}=240$ GeV, where the dark scalar can be singly produced via its mixing with $h$: $e^{+}e^{-}\to Z^{*}\to Z\,\phi$, and the dark photons are pair produced subsequently by the prompt decay of $\phi$: $\phi\to\gamma^{\prime}\gamma^{\prime}$. To our knowledge, searches for long-lived dark photons in this signal process have not been studied [^1] See Ref. Fuks:2025jrn, however, for a similar study for the proton-proton colliders.. For other phenomenological studies on theoretical scenarios involving both a dark Higgs boson and a dark photon, we refer to, e.g., Refs. Wells:2008xg; Ahlers:2008qc; Gopalakrishna:2008dv; Batell:2009yf; Weihs:2011wp; Davoudiasl:2013aya; Curtin:2013fra; Falkowski:2014ffa; Curtin:2014cca; Bakhet:2015pqa; Izaguirre:2018atq; Jodlowski:2019ycu; Araki:2020wkq; Foguel:2022unm; Cheung:2024oxh; Araki:2024uad.
@@
We confine ourselves to the mass ranges $m_{\phi}\in[25~\text{GeV},140~\text{GeV}]$ and $m_{\gamma^{\prime}}\in[2~\text{GeV},m_{\phi}/2]$. For these mass ranges, the current bounds on the mixing parameters $\epsilon$ and $\theta$ stem primarily from collider experiments. For the dark photon, the leading bounds on $\epsilon$ were obtained at the LHC collaborations CMS CMS:2019buh and LHCb LHCb:2019vmc, as well as the $B$-factory experiment BaBar BaBar:2014zli, roughly all in the order of $10^{-3}$.
@@
For $m_{\phi}$ between 100 GeV and 1 TeV, constraints around $\theta\lesssim\mathcal{O}(10^{-1})$ were achieved by searches at CMS CMS:2015hra; CMS:2018amk; CMS:2021yci and ATLAS ATLAS:2018sbw; ATLAS:2020fry; ATLAS:2020tlo; ATLAS:2022xzm. For $m_{\phi}$ between 10 GeV and 100 GeV, searches at CMS CMS:2018zvv, ATLAS ATLAS:2021hbr, LEP2 LEPWorkingGroupforHiggsbosonsearches:2003ing, and L3 Acciarri:1996aaa have placed upper bounds $\theta\lesssim\mathcal{O}(10^{-2})$. See Ref. Ferber:2023iso for a detailed summary of these existing bounds. In this work, we will simply assume $\theta=10^{-2}$ in numerical studies.
@@
For existing studies on the search prospect of dark photons and dark scalars at CEPC and FCC-ee, we refer to, for instance, Refs. He:2017zzr; He:2017ord; Biswas:2017lyg; Park:2023ygi; Liu:2017lpo; Inan:2021dir; Jiang:2019hfn; Chang:2018pjp and Refs. Liu:2017lpo; Chang:2018pjp; Cheung:2019qdr; Alipour-Fard:2018lsf; Wang:2019xvx; Ripellino:2024iem; Bhattacherjee:2025dlu; Olgoso:2025jot, respectively.
@@ II Theoretical model
We consider an extension of the SM that introduces an additional dark-sector gauge symmetry $U(1)_{D}$, under which all SM fields are uncharged. The new model includes both a dark photon $\gamma^{\prime}$ and a dark scalar $\phi$. After spontaneous symmetry breaking of the $U(1)_{D}$ via the vacuum expectation value (VEV) of the associated dark-scalar field, the dark photon acquires a mass $m_{\gamma^{\prime}}$ and an interaction between $\phi$ and two dark photons arises. The strength of this interaction scales with $m_{\gamma^{\prime}}$ and the $U(1)_{D}$ gauge coupling $g^{\prime}$. After the electroweak symmetry breaking, the dark photon will mix with the SM photon via a gauge kinetic-mixing term linking the SM hypercharge and the new $U(1)_{D}$ field strength tensors, $\epsilon F_{\mu\nu}F^{{}^{\prime}\mu\nu}$. In addition, the dark scalar $\phi$ mixes with the SM Higgs boson and thus couples to pairs of SM fermions, though the coupling strength is suppressed by the mixing parameter $\theta$. Thus, the low-energy phenomenological interaction Lagrangian employed in this study takes the form Araki:2020wkq:
@@
At the future Higgs factories, the SM Higgs bosons are dominantly produced via the Higgs-strahlung process, $e^{+}e^{-}\to Z\,h$. The dark scalar is produced in a very similar process, $e^{+}e^{-}\to Z\,\phi$, which is allowed as long as $\phi$ is light enough and it mixes with $h$. The cross section of the $e^{+}e^{-}\to Z\,\phi$ process is thus suppressed by $\theta^{2}$, and reads Djouadi:2005gi
@@
Figure 1: The production cross section of $\phi$ via the process $e^{+}e^{-}\to Z\,\phi$ rescaled with $1/\theta^{2}$ as a function of the dark-scalar mass $m_{\phi}$, computed with MadGraph5 Alwall:2014hca, at the COM energy $\sqrt{s}=240$ GeV.
In Fig. 1 we plot the scattering cross section $\sigma(e^{+}e^{-}\to Z\,\phi)$ rescaled with $1/\theta^{2}$ at the COM energy $\sqrt{s}=240$ GeV, as a function of $m_{\phi}$, computed with MadGraph5 3.4.2 Alwall:2014hca. We find these results in agreement with Eq. (2).
As mentioned earlier, in this work, we focus on $m_{\phi}\sim 10\text{--}100$ GeV. With Eq. (1), the partial decay widths of the dark scalar into a pair of dark photons and a pair of lighter SM fermions can be calculated with Araki:2020wkq
@@
The dark photon decays into a pair of SM charged fermions, with the partial decay widths Bauer:2018onh; Batell:2009yf; Araki:2020wkq; DOnofrio:2019dcp; Fabbrichesi:2020wbt; Cheung:2024oxh
@@ III The experiments and search strategies
Figure 5: A profile sketch of the inner tracker at CEPC and FCC-ee, extracted from Ref. Cheung:2019qdr.
The main detectors of CEPC and FCC-ee share similar components and geometrical structures, while their exact dimensions differ to some extent. In particular, we will focus on the inner tracker (IT) component of the main detectors in our search strategies. We thus show a profile sketch of the IT at the CEPC or FCC-ee main detector in Fig. 5, reproduced from Ref. Cheung:2019qdr. In Fig. 5, we denote a long-lived particle $i$ with polar angle $\theta_{i}$ with a dashed line. $R_{I}$ ($R_{O}$) is the inner (outer) radius of the IT, and $L_{d}$ is the half length of the IT.
@@
Table 1: Parameters of the inner tracker at CEPC and FCC-ee, extracted from Ref. CEPCStudyGroup:2018ghi and Ref. FCC:2018evy, respectively.
Further, we list in Table 1 the geometrical parameters of the inner trackers at CEPC CEPCStudyGroup:2018ghi and FCC-ee FCC:2018evy. Here, $V$ denotes the volume.
@@ IV Simulation and calculation
We perform MC simulations with the tool Pythia8.3 Bierlich:2022pfr, scanning over $m_{\gamma^{\prime}}$ and $\tau_{\gamma^{\prime}}$ for representative values of $m_{\phi}$, in order to estimate the number of signal events $N_{S}$. Concretely, we select the following four values of $m_{\phi}$: 25, 50, 90, and 140 GeV, and for each value of $m_{\phi}$, we scan over $m_{\gamma^{\prime}}$ from 2 GeV to $m_{\phi}/2$. Specifically, at every scanned grid point of $(m_{\phi},m_{\gamma^{\prime}},\tau_{\gamma^{\prime}})$, we simulate ten thousand signal events. We then translate each combination of $(m_{\gamma^{\prime}},\tau_{\gamma^{\prime}})$ into $(m_{\gamma^{\prime}},\epsilon)$ with the help of Eq. (6) and the relation $\Gamma_{\text{total}}=\displaystyle\sum_{f}\Gamma_{\gamma^{\prime}\to f\bar{f}}=\hbar/\tau_{\gamma^{\prime}}$.
In Pythia8, we turn on the process HiggsSM:ffbar2HZ at a COM energy $\sqrt{s}=240$ GeV with back-to-back electron-positron collisions. By tuning the mass of the SM Higgs boson $h$ to match that of the dark scalar $\phi$, we mimic the kinematics of the signal process $e^{+}e^{-}\to Z\,\phi$. The dark scalar $\phi$ is then set to decay promptly and exclusively into a pair of long-lived dark photons. The kinematics of the dark photons in each signal event are then used as input for calculating the acceptance $\mathcal{A}_{i}$ of each search strategy; see Eq. (7)–Eq. (10).
@@
with $i=1,2$ for the two search strategies. $\sigma(e^{+}e^{-}\to Z\,\phi)$ is computed with MadGraph5; see Fig. 1 and the relevant discussion in Sec. II. Here, $\mathcal{L}_{\text{int.}}$ denotes the integrated luminosity, with $\mathcal{L}_{\text{int.}}=22$ ab-1 at CEPC Ai:2025cpj and $\mathcal{L}_{\text{int.}}=5.6$ ab-1 at FCC-ee FCC:2018evy.
For simplicity, we assume a detector efficiency of 100% in the calculation. Additionally, the reconstruction efficiency of DVs, in principle, degrades with increasing distances from the IP. In particular, at positions close to the outer edge of the tracker we expect next to negligible DV reconstruction efficiencies. In expectation of future dedicated improvement on the experimental techniques, hardware, and search strategies, we choose to neglect such degrading effects and take an optimistic flat efficiency of 100% for the DV reconstruction.
@@ V Sensitivity reach
Figure 6: Sensitivity results of the first (upper panel) and second (lower panel) strategies. Here, the COM energy is set at 240240 GeV and the scalar-Higgs mixing angle is fixed at $10^{-2}$. The gray-, green-, and pink-filled regions have been excluded by searches at LHCb LHCb:2019vmc, CMS CMS:2019buh, and BaBar BaBar:2014zli, respectively.
@@
The existing bounds from searches at LHCb LHCb:2019vmc, CMS CMS:2019buh, and BaBar BaBar:2014zli are shown in the figures as color-filled areas.
@@
In addition, we comment that for $\theta$ values smaller than $10^{-2}$, worse sensitivity reach is expected, because fewer dark scalars and thus fewer dark photons would be produced. Quantitatively speaking, if $\theta$ is reduced by a factor of, say, 10, the total dark-photon production rates would decrease by 100. For the first search strategy, the lower sensitivity reach to $\epsilon$ would be weakened by a factor of 10 while that for the second search would be impaired by about a factor of $\sqrt{10}$.
[…]